Janelle makes monthly deposits of 90 dollars into an account that pays an effective rate of interest of 6 percent. How much will she have in the account immediately after the 24th deposit? (Hint: What is the formula for an annuity?)
Answer #1

Effective Annual Interest = 6 %

Let the monthly rate of interest be R

Therefore, (1+0.06) = (1+R)^(12)

R = (1.06)^(1/2) - 1 = .004868 or 0.4868 %

Further, let us assume that all deposits by Janelle are made at the end of each month.

Therefore, total value of deposit after the 24th one = 90 x (1.004868)^(23) + 90 x (1.004868)^(22) + ...........+ 90 x (1.004868) + 90 = [{(1.004868)^(24) - 1} / {1.004868 - 1}] x 90 = $ 2285.35